OpenStudy (abdela25):

for what value of k are the graphs of y=3x+4 and 2y=kx+9 parallel? REWARDING MEDAL

OpenStudy (anonymous):

y=3x+4 y=kx/2 + 4.5

OpenStudy (solomonzelman):

y=3x+4 and 2y=kx+9 don't worry about the y-intercepts, focus on slopes y=3x 2y=kx ---> y= (1/2)k x so 3=(1/2)k k=1.5

OpenStudy (anonymous):

Sry i misread the question Parallel graphs have the same slope Make sure that the slope of kx/2 is equal to 3

OpenStudy (anonymous):

Meaning that shouldn't the answer be 6? 6x/2 = 3x

OpenStudy (solomonzelman):

y=3x+4 and 2y=kx+9 dlopes are 3x and kx91/2) they have to be equal 3x=(1/2)kx 3=(1/2)k k=1.5

OpenStudy (anonymous):

The slopes y=3x+4 and y=6x/2 + 4.5 are parallel Meaning that k is equal to 6

OpenStudy (abdela25):

i am really confused @SolomonZelman has a different answer than what @KirbyLegs has what should the answer be and how is it done

OpenStudy (anonymous):

Soloman is correct except that he forgot to put the equation into y=mx+b form. That messed up his problem. If you graph the two graphs you will see that my two equations are parallel because of their equal slope.

OpenStudy (abdela25):

so can you do it again but this time step by step with me please @KirbyLegs

OpenStudy (anonymous):

You have one graph, which is y=3x+4 You then have to put the second graph into y=mx+b form 2y=kx+9 Divide the two to each side, giving you y=kx/2 + 4.5 Are you with me so far?

OpenStudy (abdela25):

yeah i understand

OpenStudy (abdela25):

keep going

OpenStudy (abdela25):

@KirbyLegs

OpenStudy (anonymous):

Sorry I got to go. Just very quickly though Parallel graphs have the same exact slope In the one graph you have the slope is = to 3 Make kx/2 = 3 Multiply 2 to both sides kx = 3 k=3

OpenStudy (abdela25):

oh okay i get it thanks @KirbyLegs

OpenStudy (abdela25):

@SolomonZelman what do you think...i see that he did a pretty good job

OpenStudy (solomonzelman):

I divided instead of multiplying, my bad.

OpenStudy (abdela25):

so hes right

OpenStudy (anonymous):

Both lines should have same slope to be parallel. In an equation x=mx+b, the slope is m, so for y=3x+4, the slope is 3, and for 2y=kx+9, the slope is k/2, so for them to have the same slope, k/2 =3, so k = 6. !

OpenStudy (solomonzelman):

yes he is, it's 6.